OpenStudy (mathisfun13):

A right triangle has two sides that are each 12 ft long. To the nearest foot, what is the length of the third side?

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 12^2 + 12^2 = c^2 144 + 144 = c^2 ... ... ... c = ??

OpenStudy (dan815):